OpenStudy (anonymous):

A ball is thrown upward with an initial velocity of 27 m/s. How long is the ball in the air?

OpenStudy (irishboy123):

use equation of motion: \[v = u + at\] for the upward journey \[a = -9.8 \ m/s^2\] \[u = 27 \ m/s\] \[v = 0 \ m/s\] solve for t, that's how long it takes to get to top of its upwards journey but you want 2 x t as it has to drop back again